# Task Facet Learning: A Structured Approach to Prompt Optimization

### v1

@@ 3 UniPrompt: Generating a Prompt that Captures Multiple Facets of a Task
1 Cluster train set, initialize history and validation accuracy arrays: $C\leftarrow$ cluster($D_{t}$), $H\leftarrow\{\}$, $V\leftarrow$ [];
2 Initialize a beam of size 2 with the initial prompt: $p_{1}\leftarrow p_{0}$ and $p_{2}\leftarrow p_{0}$ for each cluster c in C do
3       $B\leftarrow$ batches($C$);
4       for each batch $b\in B$ do
5             $M\leftarrow$ mini-batches($B$)
@@ 8                   Evaluate the best prompt on mini-batch: $a_{m}\leftarrow$ LLM($m,p_{1}$)
9                   Get feedback from the expert given history of mini-batch, accuracy and task description: $f\leftarrow$ Feedback($T$, $a_{m}$, $H[m]$)
@@ 12            Combine feedbacks over a batch: $F_{b}\leftarrow$ Combine($F$)
13             Apply feedback to get updated prompts; $q_{1}\leftarrow$ apply($F_{b}$, $p_{1}$); $q_{2}\leftarrow$ apply($F_{b}$, $p_{2}$)
14             Update the beam: if not($p_{1}$ = $p_{0}$) then
15                   $p_{2}\leftarrow$ second-high-acc([$p_{1}$, $p_{2}$, $q_{1}$, $q_{2}$], b)
@@ 17      Evaluate the best prompt on validation set: $acc_{v}\leftarrow$ evaluate($p_{1}$, $D_{v}$)
19       Stop based on the validation accuracy: $c\leftarrow$ early-stop-criteria($V$)
20       if $c$ then
21             break
22
23return $p_{1}$;
@@ 19       Stop based on the validation accuracy: $c\leftarrow$ early-stop-criteria($V$)
3.1 Task facet initialization using background knowledge
While existing methods for prompt optimization [12, 22] iteratively edit a human-written prompt, our goal is to generate the entire prompt from scratch (specifically, based only on a one line description of the task). Therefore, one option is to initialize the prompt using only the task description. We call this Task Description Initialization and initialize the prompt with a single section titled Introduction containing the input task description.
In the second kind of initialization, we finetune a language model to generate a prompt with multiple diverse sections such as Introduction, Tricks, and Corner Cases; similar to the initial prompt that a human prompt engineer may produce. Here we use the world knowledge embedded in language models that are trained on extensive text data. The language model is finetuned to output the entire section contents given the task description and the section title. Specifically, we train a Llama2-13B model using a synthetic dataset generated using GPT-4. Across a wide range of tasks, we first use use GPT-4 to identify the important sections given the task, section descriptions and a few examples (See Appendix A.6 for the prompt). We then fine-tune a language model $\phi$, that is Llama2-13B, in a supervised setting on a synthetically created dataset of the form $\langle T,P_{i},S_{i},{\bf S^{\prime}}_{\text{gold}}\rangle$ where $T$ denotes the task description, $P_{i}$ is the prompt generated using the fine-tuned model sequentially before section $S_{i}$ (i.e., it would have optimized sections $\{S_{1},S_{2},...S_{i-1}\}$), and ${\bf S^{\prime}}_{\text{gold}}$ is the gold section contents. $\phi$ is finetuned on the objective: $\min_{\phi}[-\log(p_{\phi}({\bf S^{\prime}}_{\text{gold}}|T,P_{i},D_{t},S_{i}))]$. We then sample 10 candidates for each section using the fine-tuned model and evaluate each section added incrementally against $D_{v}$ and select the best performing candidate for each section. More details on data collection and training of SLM are given in the Appendix A.4. Examples of one line task description and the Llama-initialized prompt for different tasks are in Appendix A.3.
3.2 Task facet refinement using examples
Extracting task-relevant concepts from a set of examples to refine a prompt is a complex problem comprising multiple steps. Given a set of incorrect predictions, one needs to analyze what went wrong in each prediction, form hypotheses, aggregate the hypotheses to identify specific concepts that are relevant for the task, and then for each concept, one needs to attribute which facet/section of the current prompt needs to be edited to incorporate the concept. These operations are highly model-specific and are difficult to execute reliably. Therefore, for this stage, we exclusively rely on an “expert” auxiliary LLM such as GPT-4.
The whole process is executed through a series of prompts to the expert LLM. First, we prompt the expert LLM to diagnose mistakes (feedback) in each example given the answer and chain-of-thought reasoning produced by the solver LLM. Subsequently, we use this feedback to generate precise edits for the prompt that may fix the error. These individual edits are then aggregated over a mini-batch and fed back into the same LLM, which then identifies a few major edits to be applied to the current prompt. To aid in identifying major edits that correspond to generalizable facets, we propose to cluster the examples as a preprocessing step and create clustered batches, such that each cluster shares some common facet of the task.
3.2.1 Preprocessing step: Clustering examples to facilitate facet identification
We explore two approaches for clustering: topic based clustering, and feedback based clustering.
Topic-Based Clustering. Given a set of examples, we identify $l$ topics that would span the entire train set. This type of clustering is motivated by the observation that solver LLM may make similar mistakes on examples from the same topic. Hence, for such examples, a common edit to the prompt could improve predictions for all the examples. To obtain the clusters, the expert LLM is prompted (for prompt see Appendix A.7) to provide a broad sub-topic $t_{i}$ for each question. Then the resultant list of sub-topics $\{t_{1},t_{2},\dots,t_{N}\}$ is again clustered into $k$ topics $\{t^{\prime}_{1},t^{\prime}_{2},\dots,t^{\prime}_{l}\}$ by prompting the expert LLM. Based on this clustering, each example ${q_{i},a_{i}}$ is assigned a cluster $t^{\prime}_{j}$ corresponding to $t_{i}$.
Feedback-Based Clustering. Another way to find examples that share similar task facets is the feedback they receive based on the initial prompt’s predictions. For instance, for a physics-based task, if two examples from different topics obtain the same feedback from the expert LLM to edit the “Rules” section of the prompt and include the statement, “Draw all forces on each body before writing the equations", then we argue that such examples can be clustered. This type of clustering makes the broad edit identification step easier. To obtain the clusters, we first evaluate all training examples against the initial prompt $p_{0}$ and store the feedback $f_{i}$ from the expert LLM, corresponding to each incorrectly answered example ${q_{i},a_{i}}$ (all the correctly answered questions form one cluster). We then prompt the expert LLM to cluster these feedbacks $\{f_{1},f_{2},\dots,f_{N}\}$ into $l$ clusters $\{f^{\prime}_{1},f^{\prime}_{2},\dots,f^{\prime}_{l}\}$ (see Appendix A.8). For each cluster $f^{\prime}_{j}$, we create a batch of examples ${q_{i},a_{i}}$ corresponding to each $f_{i}$ in $f^{\prime}_{j}$.
3.2.2 Generating prompt edits that generalize to multiple examples
Two-tier Feedback. To encourage generalizable feedback from the expert LLM that applies to more than one example, we obtain feedback at two levels: mini-batch and batch. Given a batch (created using clustering discussed above), we break it up into mini-batches. For each mini-batch $m$, we consider the questions for which the solver LM answered incorrectly with the current prompt. For these questions, we construct a prompt consisting of the questions, the chain-of-thought produced by the solver LLM, their incorrect predicted answers and the ground-truth answers. We ask the expert to provide one feedback for the mini batch (prompt is provided in Appendix A.9). In the prompt, we specify the following types of edits that the expert LLM can suggest: add a section or subsection, delete a section or subsection, and edit a section or subsection.
Given the different feedback texts for mini-batches within a batch $b$, the next step is use another call to the expert LLM to summarize the given set of edits and provide a single summarized section update. This combination ensures some degree of smoothness at every update which helps stabilize training. To make sure that the expert is able to generate generalizable edits, we additionally provide a random set of incorrect examples that are not in the current batch and ask it to suggest an edit based on the existing edits that can correct the errors. As before, the class of edits allowed is the same.
History for effective exploration: To ensure comprehensive and non-repetitive exploration of prompts, we also provide the batch-level history of edits in the mini-batch-level prompt. Specifically, we provide the expert with the history of edits $H[b]$ [12, 31] for the batch $b$ and the corresponding accuracy difference that each edit caused over the previous edit for the given batch. History is presented in the format $\{e_{i},acc_{i}-acc_{i-1}\}$ where $e_{i}$ is the edit proposed at the $i^{th}$ update and $acc_{i}$ is the accuracy of the $i^{th}$ updated prompt (See Appendix A.9 for the full prompt).
@@ 3.2.3 Editing the prompt
Once the final set is edits is received for a batch, we use the expert LLM to apply edits to the current prompt (see Appendix A.10 for the prompt). An edit is accepted only if it increases the validation accuracy compared to the current prompt. We call this approach “Greedy”. Alternatively, we maintain a beam of $l$ best performing prompts based on validation accuracy, apply the edit to all the $l$ prompts, and update the beam to retain the top $l$ performing prompts. We call this method “Beam”. To avoid overfitting on the train examples (or keep adding unnecessary information to the prompt), we employ early stopping in the optimization process (More details given in Section 4).
3.3 Efficiency considerations: Computational Complexity
The computational complexity of UniPrompt is dominated by the training phase, and in particular, by the queries made to the solver and the expert LLMs; at inference time, we need a single query to the solver LLM using the learnt prompt and the given test example. This is particularly well-suited for settings such as recommendation systems, where inference time efficiency is critical. In contrast, methods such as MedPrompt [21] require multiple calls at test time for a single test example. Let us now consider the compute complexity in terms of the number of expert or solver LLM calls made per epoch, stage-wise.
Clustering: First, we evaluate all the training examples using the current prompt. Second, for every wrongly predicted example, we obtain feedback from the expert LLM. Third, for the given set of feedbacks, we use a single call to cluster it into $l$ clusters. Each of the above steps incurs $O(N)$ queries, so the total query complexity of the clustering stage is $O(N)$. Finally, for each example, i.e., (question, answer) pair, we simply map it to the $l$ clusters (no LLM calls). This is a one-time cost.
Mini-batch feedback and Batch-level aggregation: At a given epoch, we evaluate every question in the mini-batch using the current prompt and the solver LLM ($N$ queries overall). Next, we obtain one feedback over all the wrong questions in the mini-batch $m$ ($N/|m|$ queries). We use one call to aggregate these feedbacks. For prompt selection, we evaluate $2k$ prompts on the batch $b$, where $k$ is the beam size ($O(2k|b|)$ queries per batch). Hence overall query complexity is $N+N/|m|+2kN+1$ or $O(kN)$. In practice, $k$ is a small number (2 in our experiments).
Assuming LLM throughput of 0.5 queries per second, a training + validation set of 300 examples, $l=10$ clusters, beam size $k=2$ and 20 epochs, it takes under 7 hours to train.
4 Experiments
Datasets We perform evaluation on four standard datasets : (1) Ethos [20], (2) ARC [4], (3) MedQA [13], and (4) GSM8K [5]. Ethos, ARC, and MedQA contain multiple choice questions, and GSM8K contains questions with integer answers.
@@ 4 Experiments
Implementation details We set the initial prompt $p_{0}$ for each task as the one line task description and use 200 data points $(q_{i},a_{i})$ as the train set. We use GPT-3.5-Turbo as the solver model, that is used in a chain-of-thought setting. For Feedback and Combine, we use GPT-4 as the expert. We maintain a beam size of 2. Mini-batch sizes (and batch sizes) are constrained by the context length of GPT-4. We find that mini-batch sizes 3 to 5 and batch sizes 5 to 7 work the best for our datasets. The temperature of the LLMs for our method is set to 0 for reproducibility of results. We employ early stopping at batch-level in UniPrompt. We find that a $patience$ value of 5 works best for the given tasks.
@@
Baselines We compare UniPrompt with the following techniques: (1) Task Description: prompt is the one line task description that we use to initialize UniPrompt; (2) Chain-Of-Thought (or CoT) prompting [14]; (3) Expert Prompt: the prompt optimized by humans taken from prior works [21]; (4) OPRO [31], that uses LLMs for discrete optimization over text prompts; (5) ProTeGi [22] that proposes textual gradients and selects edits to prompts using bandit techniques; (6) Evoke [12] that uses two instances of LLM, one that scores the current prompt, and the other that edits the prompt; (7) EvoPrompt [7] that uses genetic algorithms to search through the space of prompts; and (8) MedPrompt [21] that effectively combines multiple prompting strategies such as kNN-based few-shot selection, CoT reasoning, and ensembling.
Metrics For each dataset, we perform 5-fold validation (we split the train set into 5 folds) and report the final test accuracy as the average accuracy of the 5 runs on the test sets.
4.1 How does UniPrompt compare to existing methods in zero-shot setting
@@
In the zero-shot setting, we do not include labeled examples in the prompt for any of the compared methods. We report results for four versions of our method in Table 1; they differ in the initial prompt—task description, expert prompt, or the prompt generated by the fine-tuned Llama model (discussed in Section 3.1)—and the combining strategy (from Section 3.2.3)—beam search vs greedy.
@@
UniPrompt variants significantly outperform the baselines including CoT and the state-of-the-art prompt optimization techniques like ProTeGi that crucially leverage LLMs for performing iterative prompt edits. We achieve maximum gains on the Ethos dataset with a $18.2\%$ increase in accuracy over the expert prompt. Further, we see accuracy increases of $4.0\%$ on MedQA, $3.5\%$ on GSM8k, and $7.6\%$ on ARC-Challange datasets.
An example of evolution of prompts using our algorithm is given in Appendix4. It starts with a simple description of task and adds important details like differentiating between hate speech and rudeness. In contrast, ProTeGi [22] yields a rather terse prompt on the same dataset: “Does the following text contain language that targets a group of people based on their religion, gender, or other personal characteristics?”. This prompt does not provide various facets of the task, the way our method infers or an expert would typically write.
@@
Table 1: Test accuracies for the compared methods with GPT-3.5-Turbo as the solver model in the zero-shot setting (best in bold; second best underlined). The last four rows correspond to our proposed method.
@@
UniPrompt (Init = Expert Prompt) + Beam 84.0 86.0 52.3 82.4
@@
The training curves in Figure 3 show that our method initially performs edits on the prompt that simultaneously increase the train as well as the validation accuracy. After about 10 or 15 iterations (each batch update is an iteration), validation accuracy decreases while train accuracy continues increasing, indicating overfitting; which we overcome by using early stopping.
4.2 How does UniPrompt perform compared to SOTA in the few-shot setting?
@@
We now compare the performance of UniPrompt, in the few-shot setting with GPT-4 as the solver model, to the SOTA prompting technique MedPrompt [21] that employs three key components: (1) few-shot prompting, where they dynamically select five relevant examples using k-nearest neighbors (kNN) clustering in the embedding space; (2) Chain-of-Thought (CoT) reasoning on the selected examples; and (3) self-consistency and ensembling with option shuffling. From Table 2, we observe that UniPrompt (first col), which requires only one call at inference time, performs almost as well as MedPrompt (last col), which requires five calls, on the MedQA, PubMedQA, and MedMCQA datasets. As we incrementally add components such as kNN few-shot, CoT, and Ensemble to our prompt, we see a significant increase in accuracy of $4.35\%$ on average across all datasets.
4.3 Results on a Real-world Task: Search Query Intent

### v2

@@ 3.1.2 Obtaining generalizable prompt edits
1 $C\leftarrow$ cluster($D_{t},p_{0}$), initialize history $H\leftarrow\{\}$, and validation accuracies $V\leftarrow$ [];
2 Initialize a beam of size 2 with the initial prompt: $p_{1}\leftarrow p_{0}$ and $p_{2}\leftarrow p_{0}$
3 for epoch $e$ and each cluster c in C do
4
5      for each batch $b\in\texttt{batches}($C)))) do
@@ 8                   Evaluate the best prompt on mini-batch: $a_{m}\leftarrow$ LLM($m,p_{1}$)
9                   Get expert feedback: $f\leftarrow$ Feedback($T$, $a_{m}$, $H[m]$)
@@ 12            Combine feedbacks over a batch: $F_{b}\leftarrow$ Combine($F$)
13             Apply feedback to get updated prompts: $q_{1}\leftarrow$ apply($F_{b}$, $p_{1}$); $q_{2}\leftarrow$ apply($F_{b}$, $p_{2}$)
14             Update the beam: if not($p_{1}$ = $p_{0}$) then $p_{2}\leftarrow$ second-high-acc([$p_{1}$, $p_{2}$, $q_{1}$, $q_{2}$], b)
@@ 16      Evaluate the best prompt on validation set: $acc_{v}\leftarrow$ evaluate($p_{1}$, $D_{v}$)
18       if early-stop-criteria ($V$) then break
19       if recluster($e$) then $C\leftarrow$ cluster($D_{t},p_{1}$)
20return $p_{1}$ as $P^{*}$;
@@ 3.1.3 Editing the prompt
Once the final set of edits is received for a batch, we use the expert LLM to apply edits to the current prompt (See Appendix A.13 for the prompt). An edit is accepted only if it increases the validation accuracy (Greedy). Alternatively, we maintain a beam of 2 best performing prompts based on validation accuracy, apply the edit to the two prompts, and update the beam to retain the top 2 performing prompts (Beam). To avoid overfitting on the train examples (or adding unnecessary information to the prompt), we employ early stopping in the optimization process (more details in Section 4).
3.2 Prompt Initialization
We use two types of initialization: (1) task description, i.e., $p_{0}$ has a single section titled Introduction containing the input task description. (2) finetune Llama2-13B model to generate a prompt with sections such as Introduction, Tricks, and Corner Cases, similar to the initial prompt that a human prompt engineer may produce. To finetune, we use GPT-4 generated data consisting of (task description, section title, section contents) triples. Details and examples are in Appendices A.4 and A.7.
Computational Complexity:
The complexity of clustering and of getting mini-batch and batch-level feedbacks per epoch is $O(N)$ expert LLM queries, where $N$ is the number of training examples. Details are in Appendix A.3.
4 Experiments Setup
Datasets: We perform comprehensive evaluation on five standard datasets: (1) Ethos (Mollas et al., 2020), (2) ARC (Clark et al., 2018), (3) MedQA (Jin et al., 2021), (4) GSM8K (Cobbe et al., 2021), and (5) BBH (Suzgun et al., 2022). Ethos, ARC, and MedQA contain multiple choice questions, and GSM8K contains questions with integer answers. BBH is a subset of 10 tasks, spanning 4 main categories, from the challenging BIG-Bench benchmark that requires multi-step reasoning. In addition, we also evaluate UniPrompt on the medical QnA datasets used in the MedPrompt (Nori et al., 2023) work; as well as two popular code generation datasets, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
@@ 4 Experiments Setup
Implementation details: We set the initial prompt $p_{0}$ for each task as the one-line task description. We use 200 examples as the train set, 200 examples as the test set, and 100 examples as the validation set for all the compared methods. We use GPT-3.5-Turbo as the solver model. For Feedback and Combine in UniPrompt, we use GPT-4 as the expert (see ablation in Section 5.5). We maintain a beam size of 2. Mini-batch sizes (and batch sizes) are constrained by the context length of GPT-4. We find that mini-batch sizes 3 to 5 and batch sizes 5 to 7 work the best for our datasets. The temperature of the LLMs for our method is set to 0 for reproducibility. We employ early stopping at batch-level in UniPrompt.
@@
Baselines: We compare UniPrompt with the following techniques and baselines: (1) Task Description: prompt is the one line task description that we use to initialize UniPrompt; (2) Chain-Of-Thought (or CoT) prompting (Kojima et al., 2024); (3) Expert Prompt: the prompt optimized by humans taken from prior works (Nori et al., 2023); (4) OPRO (Yang et al., 2023), that uses LLMs for discrete optimization over text prompts; (5) ProTeGi (Pryzant et al., 2023) that proposes textual gradients and selects edits to prompts using bandit techniques; (6) Evoke (Hu et al., 2024) that uses two instances of LLM, one that scores the current prompt, and the other that edits the prompt; (7) EvoPrompt (Guo et al., 2023) that uses genetic algorithms to search through the space of prompts; (8) TextGrad (Hou et al., 2023), state-of-the-art framework for automatic differentiation of prompts via text; (9) DSPy (Khattab et al., 2024), a recent programming model for optimizing LLM prompts; and (10) MedPrompt (Nori et al., 2023), a state-of-the-art prompt composition method.
5 Results and Analysis
We present detailed quantitative and qualitative results, along with key ablations.
5.1 Performance of UniPrompt
@@
We start with the zero-shot setting, where we do not include labeled examples in the prompt for any of the compared methods. We report results for two versions of our method in Table 1, which differ in the combining strategy (from Section 3.1.3)—beam search vs greedy.
@@
UniPrompt variants significantly outperform the baselines including CoT and the state-of-the-art prompt optimization techniques like ProTeGi that crucially leverage LLMs for performing iterative prompt edits. UniPrompt is the best performing method on three out of four datasets. It achieves maximum gains on the Ethos dataset with a $18.2\%$ increase in accuracy over the expert prompt. Further, we see accuracy increases of $4.0\%$ on MedQA, $3.5\%$ on GSM8k, and $7.6\%$ on ARC-Challenge datasets. We show UniPrompt training behavior in Appendix A.19.
We also present comparisons to state-of-the-art DSPy method in the few-shot setting (8 bootstrapped_demos) using two optimization settings provided by their framework. The last two rows of Table 1 show that UniPrompt in the zero-shot setting convincingly outperforms DSPy in the few-shot setting, on three out of four datasets.
@@
Table 1: Test accuracies (%) of the compared methods with GPT-3.5-Turbo as the solver model in the zero-shot setting (best in bold; second best underlined). The two UniPrompt rows are our proposed method. We compare with few-shot methods in the last two rows (DSPy variants); *best in bold to distinguish the few-shot setting.
@@
DSPy (MIPRO v2, zero-shot) 79.7 82.8 61.9 77.3
TextGrad 79.5 76.5 50.6 81.6
@@
DSPy (BootstrapFewShotWithRandomSearch) 86.6 87.5 *68.5 74.3
DSPy (MIPRO v2, few-shot) 84.0 86.0 62.9 79.7
Qualitative Analysis: ProTeGi and TextGrad also adopt batching by randomly sampling from training examples where the solver LLM made mistakes. In the early iterations of optimization, there can be many such examples. So, do our key observations and hypotheses (beginning of Section 3) hold empirically? We give some evidence below.
1. Employing clustering to create batches (Section 3.1): An example feedback obtained on the Ethos dataset using UniPrompt is shown below:
The instruction should include..potential harm or violence implied, as well as any discriminatory or derogatory language used…towards a particular religious group.
The instruction should include …think about the impact of the statement on the targeted individual or group.
The instruction should…language that prompts hatred or discrimination towards a particular gender.
To contrast, we employ random batching as in the standard prompt optimization techniques, on the same dataset. The feedback obtained, given below, is relevant for the task, but fails to identify specific concepts.
The instruction should include a clear definition of hate speech…
The instruction should include examples of hate speech, guidance on identifying hate speech…
The former feedback (UniPrompt) captures the facet of measuring impact on the targeted entity whereas the latter only captures religious and harm-based aspect of hate speech. The same expert LLM is able to identify different facets due to clustered batches.
2. Employing two-tier feedback: In Section 3, we argued that employing two-tier feedback strategy to aggregate the feedback texts encourages the expert LLM to propose edits that are generalizable. The following feedback is received on the Ethos dataset after the aggregation:
The instruction should…consider whether the statement contains discriminatory, derogatory, or violent language that promotes hatred or harm towards a particular group, such as based on religion and gender.
We see that two-tier feedback helps in distilling important aspects of the task implicit in the examples, rather than directly using or rephrasing the (limited) examples.
Results on BBH: From Table 2, it is evident that UniPrompt shows a significant improvement over OPRO (that also evaluates on these tasks in their paper) for a majority of tasks. It achieves significantly higher accuracy in Boolean Expression (92.37% vs. 78.74%), Date Understanding (81.96% vs. 52.59%), and Navigate (77.16% vs. 51.74%).
Table 2: Test accuracies (%) on BBH dataset with GPT-3.5-Turbo as the solver model
Task Init OPRO UniPrompt
Algo & Multi-Step Arithmetic Reasoning
Bool Exp. 83.64 78.74 92.37
Logical Ded. 29.53 38.97 39.62
Navigate 60.95 51.74 77.16
Natural Language Understanding
Snarks 67.00 67.88 74.30
Disamb. QA 53.30 57.43 67.05
Fallacies 57.60 53.14 57.90
Use of World Knowledge
Causal Judg. 54.29 57.24 59.37
Movie Rec. 58.04 77.81 71.80
Dates 74.21 52.59 81.96
Multilingual Knowledge & Reasoning
Salient Trans. 42.59 50.61 50.77
5.2 Comparison with MedPrompt
@@
MedPrompt (Nori et al., 2023) is a recent, competitive prompting technique without any training component. It employs three key ingredients: (1) few-shot prompting, where five relevant examples are selected using k-nearest neighbors (kNN); (2) CoT reasoning on the selected examples; and (3) self-consistency and ensembling with option shuffling at inference time. They evaluate on 4 medical datasets (that none of the competing methods in Table 1 evaluate on) using GPT-4 as the solver model. So, we compare UniPrompt in the same setting in Table 8 (in Appendix A.8). UniPrompt (first row), which requires only one call at inference time, performs almost as well as MedPrompt (last row), which requires five calls, on three out of four datasets. As we incrementally add kNN few-shot, CoT, and ensembling to our prompt, we see a significant increase in accuracy of $4.35\%$ on average across all datasets.
5.3 Performance on generation tasks
Our evaluations so far have been on multiple-choice QnA, math, and classification datasets. We now evaluate UniPrompt on generating code given a natural language specification. We use HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) datasets consisting of Python coding problems. We initialize with a simple prompt, “You are a software engineer. You are given a function signature and a description of the function. You have to complete the function.” We use GPT-4-Turbo as both the solver and the expert LLM.
HumanEval does not have train or validation sets. So, we take random 100 examples from MBPP as train. Similarly, for MBPP, we take random 50 examples from HumanEval as train. We evaluate the final prompts on HumanEval and MBPP test sets. The results are given in Table 4 (in Appendix A.8). The metric is % of solved coding problems (evaluated using the provided test cases) in the datasets. The prompts produced by UniPrompt outperform standard prompting of LLMs.
5.4 Results on a Real-world Task